# On the possible contributions of two nearby blazars to the NGC 4151 neutrino hotspot

### v1

@@ Abstract
Results. Our models predict neutrino emission peaking at $\sim$1017 eV for both blazars, with fluxes of ${\sim}10^{-12}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$. This indicates their contribution to the $\sim$10 TeV neutrino signal observed from the direction of NGC 4151 is minor. While detection with current facilities is challenging, both sources should be detectable by future radio-based neutrino telescopes such as IceCube-Gen2’s radio array and GRAND, with 4FGL J1210.3+3928 being the more promising candidate.
@@ 1 Introduction
However, two gamma-ray loud blazars, 4FGL J1210.3+3928 and 4FGL J1211.6+3901, are located 0.08 and 0.43 respectively from NGC 4151 (Buson et al. 2023; Murase et al. 2024). While 4FGL J1210.3+3928 lies within the 68% confidence region of the neutrino excess attributed to NGC 4151, 4FGL J1211.6+3901 is located between two separated 95% confidence contours (Abbasi et al. 2024b) as shown in Fig. 1. Since blazars have been promising neutrino emitters, there might be a potential contribution from these two blazars to the observed neutrino signal attributed to NGC 4151. In this work, we investigate possible contributions of these nearby blazars to the neutrino excess close to NGC 4151 by numerical modeling of the observed multi-wavelength emission of 4FGL J1210.3+3928 and 4FGL J1211.6+3901. We discuss the detectability of the predicted neutrino flux from the two blazars by next-generation neutrino telescopes. Throughout this paper, we set the cosmological constants as $H_{0}=70$ km s-1 and $\Omega_{\textrm{m}}=0.3$.
Figure 1: The location of 4FGL J1210.3+3928 and 4FGL J1211.6+3901 with respect to NGC 4151. The black cross, dash-dotted and dotted black lines corresponds respectively to the best-fit location of the neutrino source and its 68% and 95% confidence regions (Abbasi et al. 2024b).
@@ 2 Data
For the rest of multi-wavelength data, optical data was taken from the SDSS Photometric Catalogue, Release 9 (Ahn et al. 2012). Infrared emission was measured by the Wide-field Infrared Survey Explorer (WISE) and the corresponding fluxes were taken from the WISE All-Sky Data Release (Cutri et al. 2012b, a). The source was also detected in 1.4 GHz by the FIRST Survey in 1993 (White et al. 1997).
@@
A second blazar, 4FGL J1211.6+3901, is classified as BL Lac object as well (Rector et al. 2000) with a redshift of $z=0.89$ based on spectroscopic measurements by Della Ceca et al. (2015). 4FGL J1211.6+3901 is detected in gamma rays with $\sim 5.4\sigma$ significance (Ballet et al. 2024; Abdollahi et al. 2022). The source has a gamma-ray flux of $(7\pm 2)\times 10^{-11}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ s^{-1}}$ in the energy range of 1-100 GeV and variability index of 15.52. The spectral shape follows a power law with photon index of $1.8\pm 0.2$.
@@ 3 Numerical Modeling
For each state, we numerically model multi-wavelength emission using the time-dependent code AM3 (Klinger et al. 2023), which solves a system of coupled differential equations describing the transport of particles interacting in the jet in a self-consistent way. Motivated by the possible neutrino emission, we start with a model where all radiation [^2] Radio fluxes cannot be explained within the one-zone model framework since the compact region is optically thick to low-frequency radio emission due to synchrotron self-absorption. originates from radiation processes of electrons and protons in the jet. We assume that both electrons and protons are accelerated in the source to power-law spectra [^3] Parameters with or without prime refer to the values in the jet or observer’s frame respectively. $dN/d{\gamma^{\prime}}_{\mathrm{e,p}}\propto{\gamma^{\prime}}_{\mathrm{e,p}}^{-\alpha_{\mathrm{e,p}}}$ with spectral indices $\alpha_{\mathrm{e,p}}$, spanning a range of Lorentz factors from ${\gamma^{\prime}}_{\mathrm{e,p}}^{\mathrm{min}}$ to ${\gamma^{\prime}}_{\mathrm{e,p}}^{\mathrm{max}}$. The energy spectra of the electrons and protons are normalized to the corresponding total electron and proton luminosities, $L^{\prime}_{\mathrm{e}}$ and $L^{\prime}_{\mathrm{p}}$. These particles are then injected into a single spherical blob of size $R^{\prime}$ (in the comoving frame of the jet) moving along the jet with Lorentz factor $\Gamma$, where there is a homogeneous and isotropic magnetic field of strength $B^{\prime}$. We assume the jet is observed at an angle $\theta_{\textrm{obs}}=1/\Gamma_{\textrm{b}}$ relative to its axis, resulting in a Doppler factor of $\delta_{\textrm{D}}\approx\Gamma_{\textrm{b}}$. We account for the high-energy gamma-ray absorption due to extragalactic background light (EBL) using the model by Domínguez et al. (2011). The best-fit parameter values were found by minimizing the reduced $\chi^{2}$ with the Minuit package (James & Roos 1975). The results of the leptohadronic modeling for 4FGL J1210.3+3928 and 4FGL J1211.6+3901 are shown in Fig. 3 and Fig. 4 respectively. The values of the leptohadronic parameters can be found in Table 1.
@@
Figure 3: Leptohadronic model for 4FGL J1210.3+3928. The green solid line correspond to the multi-wavelength photon emission. The components are shown with dashed line. The green dash-dotted line corresponds to the neutrino spectrum. The grey shaded area is a neutrino flux from NGC 4151 under the assumption of power-law with spectral index $\gamma=2.7$ (Abbasi et al. 2024a). The black dash-dotted line shows 10 year sensitivity of IceCube-Gen2 optical array (Aartsen et al. 2021), black dashed – 10 year sensitivity of IceCube-Gen2 radio array (Aartsen et al. 2021), black dotted line – 10 year sensitivity of GRAND 200k (Álvarez-Muñiz et al. 2020).
Figure 4: Leptohadronic model for 4FGL J1211.6+3901. The purple solid line correspond to the multi-wavelength photon emission. The components are shown with dashed line. The purple dash-dotted line corresponds to the neutrino spectrum. The grey shaded area is a neutrino flux from NGC 4151 under the assumption of power-law with spectral index $\gamma=2.7$ (Abbasi et al. 2024a). The black dash-dotted line shows 10 year sensitivity of IceCube-Gen2 optical array (Aartsen et al. 2021), black dashed – 10 year sensitivity of IceCube-Gen2 radio array (Aartsen et al. 2021), black dotted line – 10 year sensitivity of GRAND 200k (Álvarez-Muñiz et al. 2020).
Figure 5: Leptonic models for both quiescent (blue) and flaring (red) states. The solid line correspond to the multi-wavelength photon emission. The solid grey line corresponds to the emission of the host galaxy.
@@ 4 Results
The best-fit values of the leptohadronic models for both sources have typical values for HBL sources (Rodrigues et al. 2024). The characteristic size of the emission zone is $5\times 10^{16}$ cm with a magnetic field strength of 0.05 – 0.1 G. Both sources require the presence of a high-energy electron population (with Lorentz factors ranging from ${\sim}10^{4}$ to ${\sim}10^{6}$) in order to explain the optical and X-ray fluxes. To produce neutrinos, highly relativistic protons must be present in the jet. The maximum proton energy in the obtained models is ${\sim}10^{18}$ eV for 4FGL J1210.3+3928 and ${\sim}10^{17}$ eV for 4FGL J1211.6+3901 which correspond to the energies of ultra-high-energy cosmic rays (UHECR). Similar values for the maximum proton energies were predicted for a different HBL, Mkr 421, in Dimitrakoudis et al. (2014).
@@
Our leptohadronic modeling predicts neutrino fluxes of $\sim 10^{-12}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$ peaking above $\sim$10 PeV for both blazars. These predicted fluxes suggest that the contribution of these sources to the neutrino signal observed by IceCube in the direction of NGC 4151 is minor. As shown in Figs. 3 and 4, despite the improved PeV-range sensitivity of the optical array in the next-generation neutrino observatory IceCube-Gen2 (Aartsen et al. 2021), detection of PeV neutrinos from these blazars remains challenging. Detection of neutrinos at $10^{17}$ eV, where the emission peaks, requires either the radio array of IceCube-Gen2 or a detector like GRAND (Álvarez-Muñiz et al. 2020). The predicted neutrino emission from 4FGL J1210.3+3928 should be detectable by both facilities. The neutrino flux from 4FGL J1211.6+3901 lies at the edge of GRAND 200k’s 10-year sensitivity and slightly above the 10-year sensitivity of IceCube-Gen2’s radio array, suggesting possible detection with IceCube-Gen2. We note that the accuracy of the directional reconstruction is expected to be better than 1 for the IceCube-Gen2 radio array (Aartsen et al. 2021) and better than 0.5 for GRAND (Álvarez-Muñiz et al. 2020). In case of the detection of multiple neutrinos, these values can be further improved, leading to a possible spatial separation of signals from 4FGL J1210.3+3928 and 4FGL J1211.6+3901.
@@ 5 Discussion and Conclusions
The IceCube analysis has found around 30 signal neutrinos from NGC 4151 in 10 years of data (Abbasi et al. 2024a). However, two gamma-ray bright blazars, 4FGL J1210.3+3928 and 4FGL J1211.6+3901, are located 0.08 and 0.43 from NGC 4151 respectively. Blazars are promising neutrino emitters and can contribute to the observed IceCube signal. We modeled the multi-wavelength spectrum of both blazars. The leptohadronic model of 4FGL J1210.3+3928 explained the observed electromagnetic fluxes including a GeV dip in gamma rays, which can be explained by a purely leptonic model. For 4FGL J1211.6+3901, both leptonic and leptohadronic models explain the observed data equally well. We found that the predicted neutrino flux peaks around $10^{17}$ eV for both sources, as expected for HBLs. The contribution to the observed TeV neutrino flux is expected to be subdominant.
Both 4FGL J1210.3+3928 and 4FGL J1211.6+3901 have high-energy synchrotron peaks in X-ray, indicating the presence of efficient particle acceleration. In our model, the neutrino spectrum peaks at $\sim$100 PeV due to $p\gamma$ interactions with internal synchrotron photons and the extension of cosmic-ray spectra to EeV energies. While lowering the maximum proton energy could shift the neutrino peak toward TeV energies (Dermer et al. 2014), where IceCube detected neutrinos from the direction of NGC 4151, such models would require substantially higher proton powers and fail to explain the observed broad-band electromagnetic spectra. This demonstrates the robustness of our solution within the large parameter space of one-zone radiation models, despite potential degeneracies in the model parameters (Omeliukh et al. 2024).
@@
Models for neutrino emission from the Seyfert galaxy NGC 4151 predict neutrinos at lower energies compared to our models for the two nearby blazars. While most of the models predict a cut-off of the neutrino spectrum above 10-100 TeV for Seyfert galaxies, HBLs can produce neutrinos at higher energies. Next generation neutrino telescope such as IceCube-Gen2 or GRAND may be able to solve this discrepancy by probing a higher energy range and possibly spatially discriminating two hotspots. In addition, future MeV gamma-ray missions will be able to test signatures of hadronic emission.
@@ Appendix A Model parameters
$L^{\prime}_{\mathrm{e}}$ [erg/s] $3\times 10^{40}$ $7.5\times 10^{40}$ $7\times 10^{40}$ $1.3\times 10^{41}$

### v2

# Possible contributions of two nearby blazars to the NGC 4151 neutrino hotspot

@@ Abstract
Results. The SED of 4FGL 1210.3+3928 revealed a feature that cannot be explained with a purely leptonic model suggesting the presence of protons in the jet. Our model predicts neutrino emission peaking above $\sim$1017 eV with fluxes of ${\sim}10^{-12}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$ for this source. The SED of 4FGL J1211.6+3901 can be explained with both leptonic and leptohadronic models. The contribution of these two blazars to the $\sim$10 TeV neutrino signal observed from the direction of NGC 4151 can only be minor. Still, future radio-based neutrino telescopes such as IceCube-Gen2’s radio array and GRAND may be able to detect high-energy neutrinos from these two potential neutrino sources.
@@ 1 Introduction
However, two gamma-ray loud blazars, 4FGL J1210.3+3928 and 4FGL J1211.6+3901, are located 0.08 and 0.43 respectively from NGC 4151. 4FGL J1210.3+3928 lies within the 68% confidence region of the neutrino excess attributed to NGC 4151 (Abbasi et al. 2024b) as shown in Fig. 1, making it a natural candidate for a contributing source. This source has also been associated with a neutrino hotspot in Buson et al. (2023). While 4FGL J1211.6+3901 is located outside the 95% confidence contour, we note that these contours depend on the assumed neutrino spectrum and can vary with different spectral indices (including both statistical and systematic uncertainties, Abbasi et al. 2024b, see Fig. 8 in). Given this potential variation in the contour shape and the fact that 4FGL J1210.3+3928 and 4FGL J1211.6+3901 are the only two nearby gamma-ray loud sources (Murase et al. 2024b), we include both objects in our analysis.
Since blazars have been promising neutrino emitters, there might be a potential contribution from these two blazars to the observed neutrino signal attributed to NGC 4151. In this work, we investigate the possible contributions of these nearby blazars to the neutrino excess close to NGC 4151 by numerical modeling of the observed multi-wavelength emission of 4FGL J1210.3+3928 and 4FGL J1211.6+3901. We discuss the detectability of the predicted neutrino flux from the two blazars by next-generation neutrino telescopes. Throughout this paper, we set the cosmological constants as $H_{0}=70$ km s-1 Mpc-1 and $\Omega_{\textrm{m}}=0.3$.
Figure 1: The location of 4FGL J1210.3+3928 and 4FGL J1211.6+3901 with respect to NGC 4151. The black cross, dash-dotted, and dotted black lines correspond respectively to the best-fit location of the neutrino source and its 68% and 95% confidence regions (Abbasi et al. 2024b).
@@ 2 Data
For the rest of the multi-wavelength data, optical data was taken from the SDSS Photometric Catalogue, Release 9 (Ahn et al. 2012). Infrared emission was measured by the Wide-field Infrared Survey Explorer (WISE) and the corresponding fluxes were taken from the WISE All-Sky Data Release (Cutri et al. 2012b, a). The source was also detected in 1.4 GHz by the FIRST Survey in 1993 (White et al. 1997).
@@
A second blazar, 4FGL J1211.6+3901, is classified as BL Lac object as well (Rector et al. 2000) with a redshift of $z=0.89$ based on spectroscopic measurements by Della Ceca et al. (2015). 4FGL J1211.6+3901 is detected in gamma rays with $\sim 5.4\sigma$ significance (Ballet et al. 2024; Abdollahi et al. 2022). The source has a gamma-ray flux of $(7\pm 2)\times 10^{-11}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ s^{-1}}$ in the energy range of 1-100 GeV and variability index of 15.52. The spectral shape follows a power law with a photon index of $1.8\pm 0.2$.
@@ 3 Numerical Modeling
For each state, we numerically model multi-wavelength emission using the time-dependent code AM3 (Klinger et al. 2023), which solves a system of coupled differential equations describing the transport of particles interacting in the jet in a self-consistent way. Motivated by the possible neutrino emission, we start with a model where all radiation above $\sim$10 eV originates from radiation processes of electrons and protons in the jet. We assume that both electrons and protons are accelerated in the source to power-law spectra [^2] Parameters with or without prime refer to the values in the jet or observer’s frame respectively. $dN/d{\gamma^{\prime}}_{\mathrm{e,p}}\propto{\gamma^{\prime}}_{\mathrm{e,p}}^{-\alpha_{\mathrm{e,p}}}$ with spectral indices $\alpha_{\mathrm{e,p}}$, spanning a range of Lorentz factors from ${\gamma^{\prime}}_{\mathrm{e,p}}^{\mathrm{min}}$ to ${\gamma^{\prime}}_{\mathrm{e,p}}^{\mathrm{max}}$. The energy spectra of the electrons and protons are normalized to the corresponding total electron and proton luminosities, $L^{\prime}_{\mathrm{e}}$ and $L^{\prime}_{\mathrm{p}}$. These particles are then injected into a single spherical blob of size $R^{\prime}$ (in the comoving frame of the jet) moving along the jet with Lorentz factor $\Gamma$, where there is a homogeneous and isotropic magnetic field of strength $B^{\prime}$. We assume the jet is observed at an angle $\theta_{\textrm{obs}}=1/\Gamma_{\textrm{b}}$ relative to its axis, resulting in a Doppler factor of $\delta_{\textrm{D}}\approx\Gamma_{\textrm{b}}$. We account for the high-energy gamma-ray absorption due to extragalactic background light (EBL) using the model by Domínguez et al. (2011). The best-fit parameter values were found by minimizing the reduced $\chi^{2}$ with the Minuit package (James & Roos 1975). The results of the leptohadronic modeling for 4FGL J1210.3+3928 and 4FGL J1211.6+3901 are shown in Fig. 3 and Fig. 4 respectively. We show all-flavor neutrino fluxes here. The values of the leptohadronic parameters can be found in Table 1.
@@
Figure 3: Leptohadronic model for 4FGL J1210.3+3928. The green solid line corresponds to the multi-wavelength photon emission. The components are shown with the dashed line. The green dash-dotted line corresponds to the all-flavor neutrino spectrum. The gray shaded area is a neutrino flux from NGC 4151 under the assumption of power-law with spectral index $\gamma=2.7$ (Abbasi et al. 2024a). The black dash-dotted line shows 10 year sensitivity of IceCube-Gen2 optical array (Aartsen et al. 2021), black dashed – 10 year all-flavor sensitivity of IceCube-Gen2 radio array (Aartsen et al. 2021), black dotted line – 10 year all-flavor sensitivity of GRAND 200k (Álvarez-Muñiz et al. 2020).
Figure 4: Leptohadronic model for 4FGL J1211.6+3901. The purple solid line corresponds to the multi-wavelength photon emission. The components are shown with the dashed line. The purple dash-dotted line corresponds to the all-flavor neutrino spectrum. The gray shaded area is a neutrino flux from NGC 4151 under the assumption of power-law with spectral index $\gamma=2.7$ (Abbasi et al. 2024a). The black dash-dotted line shows the 10 year sensitivity of the IceCube-Gen2 optical array (Aartsen et al. 2021), black dashed – 10 year all-flavor sensitivity of IceCube-Gen2 radio array (Aartsen et al. 2021), black dotted line – 10 year all-flavor sensitivity of GRAND 200k (Álvarez-Muñiz et al. 2020).
Figure 5: Leptonic models for both quiescent (blue) and flaring (red) states. The solid line corresponds to the multi-wavelength photon emission. The solid gray line corresponds to the emission of the host galaxy.
@@ 4 Results
The best-fit values of the leptohadronic models for both sources have typical values for HBL sources (Rodrigues et al. 2024). The characteristic size of the emission zone is $5\times 10^{16}$ cm with a magnetic field strength of 0.05 – 0.1 G. Both sources require the presence of a high-energy electron population (with Lorentz factors ranging from ${\sim}10^{4}$ to ${\sim}10^{6}$) in order to explain the optical and X-ray fluxes. To produce neutrinos, highly relativistic protons must be present in the jet. The maximum proton energy in our models reaches ${\sim}10^{18}$ eV for 4FGL J1210.3+3928, which corresponds to the ultra-high-energy cosmic ray (UHECR) regime, while it is ${\sim}10^{17}$ eV for 4FGL J1211.6+3901. Similar values for the maximum proton energies were predicted for a different HBL, Mkr 421, in Dimitrakoudis et al. (2014).
@@
In all of the presented models, the intrinsic gamma-ray luminosity, before EBL absorption, is comparable to the neutrino luminosity, maintaining energy conservation in the hadronic processes. However, in the observed spectra (as shown in Figs. 3 and 4), gamma rays with energies above TeV are strongly absorbed due to EBL, creating an apparent energetic discrepancy between gamma-ray and neutrino fluxes. We note that the electromagnetic cascade of these absorbed gamma rays could contribute to the observed MeV – GeV fluxes. However, as the intergalactic magnetic field (IGMF) has a strength $>7.1\times 10^{-16}$ G (e.g., Aharonian et al. 2023), the cascade emission can significantly dissipate. This is consistent with the lack of clear evidence for secondary cascade components in current gamma-ray observations. Given these observational constraints and the uncertainty of the IGMF strength, we did not include the intergalactic cascade component in our analysis.
Our leptohadronic modeling predicts all-flavor neutrino fluxes of $\sim 10^{-12}\leavevmode\nobreak\ \mathrm{erg\leavevmode\nobreak\ cm^{-2}\leavevmode\nobreak\ s^{-1}}$ peaking above $\sim$10 PeV for both blazars. These predicted fluxes suggest that the contribution of these sources to the neutrino signal observed by IceCube in the direction of NGC 4151 is minor. In our data analysis, we selected a high level of X-ray flux to estimate the maximum possible neutrino contribution from both blazars. The time-averaged emission state would only further reduce blazar contribution to the NGC 4151 hotspot, strengthening our conclusion about its subdominant role.
Despite the improved PeV-range sensitivity of the optical array in the next-generation neutrino observatory IceCube-Gen2 (Aartsen et al. 2021), the detection of possible neutrinos emission from 4FGL J1210.3+3928 or 4FGL J1211.6+3901 remains challenging due to their high energies. Detection of neutrinos at $10^{16}-10^{17}$ eV, where the emission peaks, requires either the radio array of IceCube-Gen2 or a detector like GRAND (Álvarez-Muñiz et al. 2020). The predicted neutrino emission from 4FGL J1210.3+3928 could be detectable by both facilities. The neutrino flux from 4FGL J1211.6+3901 lies at the edge of GRAND 200k’s 10-year sensitivity and slightly above the 10-year sensitivity of IceCube-Gen2’s radio array, suggesting possible detection with IceCube-Gen2. We note that the accuracy of the directional reconstruction is expected to be better than 1 for the IceCube-Gen2 radio array (Aartsen et al. 2021) and better than 0.5 for GRAND (Álvarez-Muñiz et al. 2020). In case of the detection of multiple neutrinos, these values can be further improved, leading to a possible spatial separation of signals from 4FGL J1210.3+3928 and 4FGL J1211.6+3901.
@@ 5 Discussion and Conclusions
The IceCube analysis has found around 30 signal neutrinos from NGC 4151 in 10 years of data (Abbasi et al. 2024a). However, two gamma-ray bright blazars, 4FGL J1210.3+3928 and 4FGL J1211.6+3901, are located 0.08 and 0.43 from NGC 4151 respectively. Blazars are promising neutrino emitters and can contribute to the observed IceCube signal. We modeled the multi-wavelength spectrum of both blazars. The leptohadronic model of 4FGL J1210.3+3928 explained the observed electromagnetic fluxes including a GeV dip in gamma rays, which can be explained by a purely leptonic model. For 4FGL J1211.6+3901, both leptonic and leptohadronic models explain the observed data equally well. We found that the predicted neutrino flux peaks around $10^{17}$ eV for both sources. The contribution to the observed TeV neutrino flux is expected to be subdominant.
Both 4FGL J1210.3+3928 and 4FGL J1211.6+3901 have high-energy synchrotron peaks in X-ray, indicating the presence of efficient particle acceleration. In our model, the neutrino spectrum peaks at $\sim$10–100 PeV due to $p\gamma$ interactions with internal synchrotron photons and the extension of cosmic-ray spectra to EeV energies. While lowering the maximum proton energy could shift the neutrino peak toward TeV energies (Dermer et al. 2014), where IceCube detected neutrinos from the direction of NGC 4151, such models would require substantially higher proton powers and fail to explain the observed broad-band electromagnetic spectra. The maximum proton Lorentz factor for the leptohadronic model of 4FGL J1211.6+3901 is substantially lower than that for 4FGL J1210.3+3928. This is caused by the fact that with the increase of proton energy, the contributions of Bethe-Heilter pairs and proton synchrotron become more significant leading to the violation of the observed X-ray spectrum and featureless gamma-ray spectrum.This demonstrates the robustness of our solution within the large parameter space of one-zone radiation models, despite potential degeneracies in the model parameters (Omeliukh et al. 2024).
@@
Models for neutrino emission from the Seyfert galaxy NGC 4151 predict neutrinos at lower energies compared to our models for the two nearby blazars. While most of the models predict a cut-off of the neutrino spectrum above 10-100 TeV for Seyfert galaxies, HBLs can produce neutrinos at higher energies. Next-generation neutrino telescopes such as IceCube-Gen2 or GRAND may solve this discrepancy by probing a higher energy range and possibly spatially discriminating two hotspots. In addition, future MeV gamma-ray missions will be able to test signatures of hadronic emission.
@@ Appendix A Model parameters
$L^{\prime}_{\mathrm{e}}$ [erg/s] $3\times 10^{40}$ $7.5\times 10^{40}$ $7\times 10^{40}$ $1.4\times 10^{41}$